They sell lottery tickets at the kiosk. A single lottery ticket wins with a probability of 0.45. With

likelihood

a) all bets win 

b) at least one lot wins 

c) exactly two lotteries win?


1
Expert's answer
2021-12-20T16:23:36-0500

Let "X=" the number of tickets which win: "X\sim Bin (n, p)"

Given "p=0.45"

a)

"P(X=n)=\dbinom{n}{n}(0.45)^n(1-0.45)^{n-n}=(0.45)^n"

b)


"P(X\geq 1)=1-P(X=0)"

"=1-\dbinom{n}{0}(0.45)^0(1-0.45)^{n-0}=1-(0.55)^n"

c)


"P(X=2)=\dbinom{n}{2}(0.45)^2(1-0.45)^{n-2}"

"=\dfrac{n(n-1)}{2}(0.45)^2(0.55)^{n-2}"